# ,, thanks: These two authors contributed equally to this work thanks: These two authors contributed equally to this work

### v1

@@ Abstract
The quantum Hall effect (QHE) is a cornerstone of topological physics, inspiring extensive explorations of its high-dimensional generalizations such as experimental realizations in synthetic systems including cold atoms, photonic lattices and metamaterials. However, the many-body effect in the higher dimensional QHE system remains poorly understood. In this study, we explore this problem by formulating the microscopic wavefunctions on a four dimensional sphere inspired by Laughlin’s seminal work. Employing a generalized pseudo-potential framework, we derive an exact microscopic Hamiltonian consisting of two-body projectors that annihilate the microscopic wavefunctions. Diagonalizations on finite system sizes show that the quasi-hole states remain zero energy while the quasi-particle states exhibit a finite gap, in consistency with an incompressible state. Furthermore, the pairing distribution is calculated to substantiate the liquid-like nature of the wavefunction. Our work provides a preliminary understanding to the fractional topological states in high dimension.
@@
The discovery of the integer quantum Hall effect (IQHE) Klitzing et al. (1980) marked a pivotal milestone in condensed matter physics research, opening up an era for studying topological phenomena. In IQHE systems, two-dimensional (2D) electrons subjected to a magnetic field form Landau levels, and the quantization of Hall conductance is intrinsically tied to Landau level wavefunction topologies Thouless et al. (1982). Moreover, many-body properties become sharp research focus following the discovery of the fractional quantum Hall effect (FQHE) Tsui et al. (1982). The ground breaking insight comes from the celebrated Laughlin wavefunction Laughlin (1983), which captures the essence of the FQHE including the fractionalization of Hall conductance and fractional excitations. Various generalizations of the Laughlin wavefunction have been proposed, including the Read-Moore state Moore and Read (1991) and the Haldane-Rezayi state Haldane and Rezayi (1988), offering new perspectives on the FQHE.
@@
The extension of Landau levels to higher-dimensional systems has been investigated in various contexts. Laudau levels on the four-dimensional (4D) sphere ($S^{4}$) have been constructed Zhang (2001) by introducing an SU(2) Yang monopole located at the origin Yang (1978a), in analogy to Haldane’s construction of Landau levels on a 2D sphere with a $U(1)$ monopole Haldane (1983). In addition, Landau levels in high dimensional flat space for both non-relativistic Li and Wu (2013); Li et al. (2013) and Dirac fermions Li et al. (2012) have been constructed. The 4D quantum Hall effect (QHE) has been applied to the study of noncommutative geometry Chen et al. (2002); Susskind (2001); Hou and Peng (2002); Zhang (2003); Hasebe (2010, 2018, 2020, 2022, 2023).Experimental advances have also demonstrated the realization of the 4D QHE in cold atom systems Bouhiron et al. (2024); Price et al. (2015), optical lattices Lohse et al. (2018); Zilberberg et al. (2018), acoustic lattices Chen et al. (2021a, b); Xu et al. (2022); Chen et al. (2023a); Lai et al. (2024), and electric circuitsPrice (2020); Wang et al. (2020); Yu et al. (2020); Zhang et al. (2020). Furthermore, the Yang monopole has been observed in cold atom simulations and metamaterials Sugawa et al. (2018); Ma et al. (2021, 2023). Nevertheless, the many-body physics underlying the 4D FQHE remains an open and intriguing challenge for researchers , which is expected to explore experimentally in synthetic dimensions Yang et al. (2024); Ozawa and Price (2019); Wu et al. (2019); Guo et al. (2021); Jiang et al. (2021); Ezawa (2021); Lustig and Segev (2021); Kanungo et al. (2022); Trautmann et al. (2024).
@@
In this article, we explore the many-body properties based on $S^{4}$ with the SU(2) monopole structure. Both the Laughlin-type many-body wavefunctions as a power of a determinant type and the Jastrow type are constructed, which are ground states of Hamiltonians with suitably designed Haldane pseudo-potentials. Exact diagonalizations are performed for the Laughlin-type wavefunction yielding the signature of an incompressible quantum liquid state. These states provide further insights for studying high dimensional topological states.
@@
We decompose the Landau level according to the IRREPS of the subgroup $SO(4)$, whose 6 generators can be constructed with t’Hooft symbol 111 The $\hat{J}_{a},\hat{K}_{a}$ satisfy commutation that $\hat{J}_{a},\hat{J}_{b}]=\text{i}\varepsilon_{abc}\hat{J}_{c},[\hat{K}_{a},\hat{K}_{b}]=\text{i}\varepsilon_{abc}\hat{K}_{c},[\hat{J}_{a},\hat{K}_{b}]=0$, see more detail in S.M.sup as
@@
Applying the branching rule to the LLL states, we arrive at $(2I,0)_{so5}=\bigoplus\limits_{j+k=I}(j,k)_{so4}$. Since $\hat{J}_{a}$ and $\hat{K}_{a}$ form two sets of inter-commutable $SU(2)$ algebra, $(j,k)_{so4}$ represents a direction product of SU(2) representations characterized by quantum numbers of $j$ and $k$, respectively.
Next we explain how to express the LLL wavefunctions by the homogeneous polynomials of the 4-component fundamental $SO(5)$ spinors $\psi_{\alpha}$ with $\alpha=1\sim 4$. Through the 2nd Hopf map, $\psi$ defines a unit vector $x_{a}=\psi^{\dagger}_{\alpha}\Gamma^{a}_{\alpha\beta}\psi_{\beta}(a=1\sim 5)$ on $S^{4}$. The $4\times 4$ Clifford algebra $\Gamma^{a}$ matrices anti-commutes each other and its concrete representation is given in S. M. Sec. II . Conversely, $\psi_{\alpha}$’s are represented by $x_{a}$ and a 2-component spinor $u=(u_{1},u_{2})^{T}$ which determines a unit vector on $S^{2}$ via the 1st Hopf map ${n}_{i}=u^{\dagger}{\sigma}_{i}u$. (see S.M. Sec. II). The LLL wavefunctions are defined as a $2I+1$-component spinor with each component the eigenstate of $I_{3}$, denoted as $f_{I_{3}}(x_{a})|II_{3}\rangle$ where $I\geq I_{3}\geq-I$. In order for a compact notation, $|II_{3}\rangle$ is mapped to a wavefunction on $S^{2}$ determined by the isospin number $I$ as $u_{1}^{I+I_{3}}u_{2}^{I-I_{3}}/\sqrt{(I+I_{3})!(I-I_{3})!}$. Then the LLL wavefunctions are expressed as
@@
For later convenience to explore the many-body wavefunctions, we first construct the 2-body states with respect to the $SO(5)$ rotational symmetry on $S^{4}$, which is significantly more complex than that case of $S^{2}$ subject to the $SO(3)$ symmetry. The product of two single particle LLL states are decomposed into different $SO(5)$ channels according to the rule that,
@@
where $0\leq n_{1}\leq 2I$, and $0\leq n_{2}\leq n_{1}$. The simplest set of the two-body states lie in the channel of $(4I-2n,0)_{so5}$ with $0\leq n\leq 2I$. An SO(5) invariant is constructed as $S(x_{a},n_{i};x^{\prime}_{a},n_{i}^{\prime})=\psi_{\alpha}(x_{a},n_{i})\mathcal{R}_{\alpha\beta}\psi_{\beta}(x^{\prime}_{a},n_{i}^{\prime})$ with $\mathcal{R}$ the charge conjugation matrix for the $SO(5)$ group satisfying the following properties, $\mathcal{R}^{2}=-1$, $\mathcal{R}^{\mathbf{T}}=\mathcal{R}^{-1}=\mathcal{R}^{\dagger}=$-$\mathcal{R}$  [^2] The explicit definition of $\mathcal{R}$ is shown in S.M. Sec. III. The $\mathcal{R}$ is defined as $\mathcal{R}=-\text{i}\Gamma_{3}\Gamma_{1}$. Under the help of $\mathcal{R}$, we derive the expressions for the 2-body states as
@@
where $\mathcal{N}_{J,M_{1};K,M_{2}}^{(4I-2n,0)}$ is the normalization coefficient that is given by $SO(5)$ reduced isoscalar (see S.M. Sec. VI), and $f_{J_{1},M_{1};J_{2},M_{2}}^{4I-2n}$ is a symmetric homogeneous polynomial with the degree of 4I2n4I-2n defined as
@@
where $C_{j_{1},m_{1};j_{1}^{\prime},m_{1}^{\prime}}^{J_{1},M_{1}}$ is the SU(2) group Celebsh-Gordon coefficient . The statistics of the 2-body wavefunciton is captured by the power of $S^{n}$: Odd and even values of $n$ correspond to the fermionic and bosonic statistics, respectively.
We move forward to the 2-body states in more complex channels with $n_{1}>n_{2}$. Let us take the case of $n_{1}-n_{2}=1$ as an example which involves the regular $SO(5)$ harmonics in the absence of Yang monopole . In analogy to the SO(5) invariant $S$, an $SO(5)$ vector, defined as $X_{a}(x_{a},n_{i};x_{a}^{\prime},n_{i}^{\prime})=\psi_{\alpha}(x_{a},n_{i})(\mathcal{R}\Gamma_{a})_{\alpha\beta}\psi_{\beta}(x_{a}^{\prime},n_{i}^{\prime})$, transforms under the IRREP $(1,1)_{so5}$ . $X_{a}$ can be decomposed into an SO(4) scalar $X_{5}$ and a 4-vector $X_{1\sim 4}$ denoted as $(\frac{1}{2};\frac{1}{2})_{so4}$(see S.M.sup .Table(S3)) $S$ and $X_{5}$ are orthogonal to each other, since the former is SO(5) invariant and the latter is a component of 5-vector. According to group theoretical analysis, the SO(5) of IRREP $(4I-2n-1,1)_{so5}$ with $0\leq n\leq 2I-1$ can be decomposed into two branches of SO(4) IRREPs as follows,
@@
Hence, we have two sets of the 2-body wavefunctions as
@@
in which $\Psi_{\frac{1}{2},\pm\frac{1}{2};\frac{1}{2},\pm\frac{1}{2}}^{(1,1)}=\frac{1}{\sqrt{2}}(X_{1}\pm iX_{2})$, and $\Psi_{\frac{1}{2},\pm\frac{1}{2};\frac{1}{2},\mp\frac{1}{2}}^{(1,1)}=\frac{1}{\sqrt{2}}(X_{3}\pm iX_{4}).$ As for the general case of $(n_{1}+n_{2},n_{1}-n_{2})_{so5}$, the 2-body wavefunctions in general IRREPs are constructed by involving high ranks of $SO(5)$ harmonics as discussed in S. M. Sec(IV).
To implement numerical calculations, we generalize the pseudo-potential formalism Haldane (1983); Jain (2007) on $S^{2}$ to $S^{4}$. The pseudo-potential is represented by a 2-body projection operator as follows,
@@
The determinant-type Laughlin wavefunction is represented as the power of the slater determinant.
@@
where $N$ is the particle number; $A_{h}$ is the abbreviation of $SO(4)$ IRREP $(j_{h},m_{h,1};k_{h},m_{h,2})$ for the $h$-th particle. The filling of Eq.(12) is $\nu=\frac{d(2I,0)}{d(2mI,0)}$ , which approaches $1/m^{3}$ in the large-$m$ limit.
@@
$\Psi^{(m=3)}_{\mbox{det},N=4}=\left(\Psi_{\mbox{det}}(x_{1,a},n_{1,i};...,;x_{4,a},n_{4,i})\right)^{3}.$ (13)
By counting the power of SO(5) spinors, the single-particle LLL orbitals involved in Eq.(13) correspond to the case of $mI=\frac{3}{2}$ with the degeneracy $d(3,0)=20$. The full shell determinant state is SO(5) invariant, hence, its power, the Laughlin determinant state, is also an $SO(5)$ singlet. Exchanging two particles results in the sign change of $(-1)^{3}$, which ensures the fermi statistics. Hence, when expanding Eq.(13), the relative wavefunction between a pair of particles $k$ and $l$ only contains
@@
Now we construct the pseudo-potential Hamiltonian for the many-body wavefunctions Eq.(12). A series of 2-body projection operators are selected to ensure that the pseudo-potential Hamiltonian is positive-definite and annihilates the wavefunction(12). Hence, such a the wavefunction is at zero energy and then a ground state. For concreteness, the special case of Eq.(13) with $m=3,N=4$ is employed as an example, i.e.. Consider two fermions in the LLL orbits with $I^{\prime}=mI=\frac{3}{2}$. The 2-particle states satisfying the fermionic statistics are decomposed into SO(5) IRREPs of $(4I^{\prime}-2r-1-s,2r+1-s)_{so5}$ with $0\leq r<I^{\prime}$ and $0\leq s\leq 2r+1$. According to the 2-body wavefunctions given in Eq.(VI.8) in S. M. Sec. (VI), the contraction number equals $2r+1$. Hence, the above IRREPs with $r=0$, $(4,0)_{so5},(5,1)_{so5}$, are orthogonal to any 2-body sector inside the wavefunction of Eq.(13) by comparing the contraction numbers. The above analysis can be easily generalize to an arbitary value of $N$ at $m=3$. Then the pseudo-potential Hamiltonian is constructed as
@@
where $P_{m,n}(k,l)$ is the projection operator of 2-particle states of the $k$-th and $l$-th ones.
@@
Next we present the exact digonalization results for the case of $m=3$ and $N=4$. The $SO(5)$ is a rank-2 Lie group with two commutable generators, which can be taken as the sum of the diagonal operators of two inter-commutable $SU(2)$ subgroups defined in Eq.(4) of every particle: $J_{3}=\sum_{l=1}^{N}J_{3}(l)$ and $K_{3}=\sum_{l=1}^{N}K_{3}(l)$. The pseudo-potential Hamiltonian is diagonalized in blocks labeled by different values of $(M_{J},M_{K})$, which are $J_{3}$ and $K_{3}$ eigenvalues, respectively. Previously analysis shows that the wavefunction of Eq.(13) is an SO(5) singlet and a ground state of the Hamiltonian Eq. (15). Diagonalizations are performed and the ground state is found non-degenerate, hence, the wavefunction of Eq.(13) is the unique ground state. Energy spectra in the blocks of $(0,0)$ and $(\frac{1}{2},\frac{1}{2})$ with spectra shown in Fig.(1). The energies of the lowest and next lowest excited states are very close at $4.80$ and $4.93$ under the unit of $\frac{\hbar^{2}}{2MR^{2}}$. The lowest excited state is also an $SO(5)$ scalar as shown in the sector of $(0,0)$, and the next lowest excited states are 35-fold degenerate forming the SO(5) IRREP of $(4,0)_{{so5}}$. Such an $SO(5)$ IRREP is decomposed into different $SO(4)$ IRREPs carrying integer quantum numbers of $(2,0)_{so4}\oplus(1,1)_{so4}\oplus(0,2)$ with dimensions of $5$, 99, and $5$, respectively, and those carrying half-integer quantum numbers of $(\frac{3}{2},\frac{1}{2})_{so4}\oplus(\frac{1}{2},\frac{3}{2})_{so4}$ with dimensions 8 and 8, respectively, as shown in Fig.(1) . More spectra details are shown on S.M.sup .Sec(VIII).
@@
We now present the spectra of excitations of the quasi-hole ($N=3$) and quasi-particle ($N=5$) sectors as shown in Fig.(2) ($a$) and ($b$), respectively. Since each SO(5) IRREP contains SO(4) IRREPS $(j_{1},j_{2})_{so4}\oplus(j_{2},j_{1})_{so4}$ symmetrically, we only present the state in the sector $(\frac{1}{2},0)$ in terms of $(M_{J},M_{K})$ as a representative. As for the quasi-hole states of $N=3$, its ground states remains at zero energy with 20-fold degeneracy, lying in the SO(5) IRREP of $(3,0)_{so5}$, which is decomposed to SO(4) IRREPs of $(\frac{3}{2},0)_{so4}\oplus(0,\frac{3}{2})_{so4}\oplus(1,\frac{1}{2})_{so4}\oplus(\frac{1}{2},1)_{so4}$. In contrast, as for the quasi-particle states with $N=5$, there exist a finite excitation gap. The first excited states form the SO(5) IRREP of $(3,0)_{so5}$, which is the same as the case of quasi-hole.
@@
Early studies of FQHE debated whether the ground state should be an incompressible quantum liquid or a Wigner crystal  Yoshioka et al. (1983); Yoshioka and Lee (1983); Yoshioka (1984). The pairing distribution function is frequently employed to reveal the incompressible nature of quantum liquids Jain (2007). We calculate the probability $h(\theta)$ of two particles separated at a distance, which is a function of the angle $\theta$ between two unit vectors from the center to particles due to the rotation symmetry. It is defined as

### v2

@@ Abstract
The quantum Hall effect (QHE) is a cornerstone of topological physics, inspiring extensive explorations of its high-dimensional generalizations such as experimental realizations in synthetic systems including cold atoms, photonic lattices and metamaterials. However, the many-body effect in the higher dimensional QHE system remains poorly understood. In this study, we explore this problem by formulating the microscopic wavefunctions on a four dimensional sphere inspired by Laughlin’s seminal work. Employing a generalized pseudo-potential framework, we derive an exact microscopic Hamiltonian consisting of two-body projectors that annihilate the microscopic wavefunctions. Diagonalizations on finite system sizes show that the quasi-hole states remain zero energy while the quasi-particle states exhibit a finite gap, in consistency with an incompressible state. Furthermore, the pair distribution is calculated to substantiate the liquid-like nature of the wavefunction. Our work provides a preliminary understanding to the fractional topological states in high dimension.
@@
The discovery of the integer quantum Hall effect (IQHE) Klitzing et al. (1980) marked a pivotal milestone in condensed matter physics research, opening up an era for studying topological phenomena. In IQHE systems, two-dimensional (2D) electrons subjected to a magnetic field form Landau levels, and the quantization of Hall conductance is intrinsically tied to Landau level wavefunction topologies Thouless et al. (1982). Moreover, many-body properties become sharp research focus following the discovery of the fractional quantum Hall effect (FQHE) Tsui et al. (1982). The groundbreaking insight comes from the celebrated Laughlin wavefunction Laughlin (1983), which captures the essence of the FQHE, including fractional Hall conductance and fractionally charged excitations. Various generalizations of the Laughlin wavefunction have been proposed, including the Read-Moore state Moore and Read (1991) and the Haldane-Rezayi state Haldane and Rezayi (1988), offering new perspectives on the FQHE.
@@
Various investigations have explored the extension of Landau levels to higher-dimensional systems Zhang (2001); Karabali and Nair (2002); Bernevig et al. (2003); Jellal (2005); Li et al. (2012); Li and Wu (2013); Li et al. (2013); Peng et al. (2025).Most notably, Landau levels on the four-dimensional (4D) sphere ($S^{4}$) have been constructed Zhang (2001) by introducing an SU(2) Yang monopole located at the origin Yang (1978a), analogous to Haldane’s construction of Landau levels on $S^{2}$ with Dirac monopole Haldane (1983). The 4D quantum Hall effect (QHE) has been applied to the study of noncommutative geometry Chen et al. (2002); Susskind (2001); Hou and Peng (2002); Zhang (2003); Hasebe (2010, 2018, 2020, 2022, 2023).Experimental advances have also demonstrated the realization of the 4D QHE in cold atom systems Bouhiron et al. (2024); Price et al. (2015), optical lattices Lohse et al. (2018); Zilberberg et al. (2018), acoustic lattices Chen et al. (2021a, b); Xu et al. (2022); Chen et al. (2023a); Lai et al. (2024), and electric circuitsPrice (2020); Wang et al. (2020); Yu et al. (2020); Zhang et al. (2020). Furthermore, the Yang monopole has been observed in cold atom simulations and metamaterials Sugawa et al. (2018); Ma et al. (2021, 2023).Nevertheless, the many-body physics of the 4D FQHE remains poorly understood, constituting an exceptionally difficult and largely unexplored frontier that may be accessible through synthetic dimensions in future experiments.Yang et al. (2024); Ozawa and Price (2019); Wu et al. (2019); Guo et al. (2021); Jiang et al. (2021); Ezawa (2021); Lustig and Segev (2021); Kanungo et al. (2022); Trautmann et al. (2024).
@@
In this article, we explore the many-body properties based on $S^{4}$ under the SU(2) monopole background. Both the Laughlin-type many-body wavefunctions as a power of a determinant type and the Jastrow type are constructed, which are ground states of Hamiltonians with suitably designed Haldane pseudo-potentials. Exact diagonalizations are performed on the corresponding Hamiltonians, revealing the signature of an incompressible quantum liquid state. These states provide further insights for studying high dimensional topological states.
@@
We decompose the LLL according to the IRREPS of the subgroup $SO(4)$, with 6 generators can be constructed with t’Hooft symbol 111 The $\hat{J}_{a},\hat{K}_{a}$ satisfy commutation that $\hat{J}_{a},\hat{J}_{b}]=\text{i}\varepsilon_{abc}\hat{J}_{c},[\hat{K}_{a},\hat{K}_{b}]=\text{i}\varepsilon_{abc}\hat{K}_{c},[\hat{J}_{a},\hat{K}_{b}]=0$, see more detail in S.M.sup (2020)
@@
Applying the branching rule to the LLL states, we obtain $(2I,0)_{so5}=\bigoplus\limits_{j+k=I}(j,k)_{so4}$. Since $\hat{J}_{a}$ and $\hat{K}_{a}$ form two mutually commuting $SU(2)$ algebras, $(j,k)_{so(4)}$ denotes the direct product of two $SU(2)$ representations characterized by quantum numbers $j$ and $k$, respectively.
Next we explain how to express the LLL wavefunctions by the homogeneous polynomials of the 4-component fundamental $SO(5)$ spinors $\psi_{\alpha}$ with $\alpha=1\sim 4$. Through the second Hopf map, one can define a unit vector $x_{a}=\psi^{\dagger}_{\alpha}\Gamma^{a}_{\alpha\beta}\psi_{\beta},(a=1,\dots,5)$ on $S^{4}$. The $4\times 4$ Clifford algebra matrices $\Gamma^{a}$ mutually anticommute and their explicit representation is presented in S.M.Sec.II. Conversely, $\psi_{\alpha}$’s can be expressed through $x_{a}$ and a two-component spinor $u=(u_{1},u_{2})^{T}$ ,defining the isospin coordinates on $S^{2}$ via the 1st Hopf map ${n}_{i}=u^{\dagger}{\sigma}_{i}u$. (see S.M.Sec. II). The LLL wavefunctions can be organized into a $(2I+1)$-component spinor, with each component being an eigenstate of $I_{3}$, denoted by $f_{I_{3}}(x_{a})|II_{3}\rangle$, where $I\geq I_{3}\geq-I$. For compactness, $|II_{3}\rangle$ is represented on $S^{2}$ by $u_{1}^{I+I_{3}}u_{2}^{I-I_{3}}/\sqrt{(I+I_{3})!(I-I_{3})!}$. Then the LLL wavefunctions are expressed as
@@
For later convenience in exploring many-body wavefunctions, we first construct the two-body states under the $SO(5)$ rotational symmetry on $S^{4}$, which is significantly more complex than the $S^{2}$ case with $SO(3)$ symmetry. The product of two single-particle LLL states is then decomposed into different $SO(5)$ channels. According to the branching rule ,
@@
where $0\leq n_{1}\leq 2I$, and $0\leq n_{2}\leq n_{1}$. The simplest set of the two-body states lie in the channel of $(4I-2n,0)_{so5}$ with $0\leq n\leq 2I$. An SO(5) invariant is constructed as $S(x_{a},n_{i};x^{\prime}_{a},n_{i}^{\prime})=\psi_{\alpha}(x_{a},n_{i})\mathcal{R}_{\alpha\beta}\psi_{\beta}(x^{\prime}_{a},n_{i}^{\prime})$ with $\mathcal{R}$ being the charge conjugation matrix for the $SO(5)$ group , satisfying $\mathcal{R}^{2}=-1$, $\mathcal{R}^{\mathbf{T}}=\mathcal{R}^{-1}=\mathcal{R}^{\dagger}=$-$\mathcal{R}$  [^2] The explicit definition of $\mathcal{R}$ is shown in S.M. Sec. III. The $\mathcal{R}$ is defined as $\mathcal{R}=-\text{i}\Gamma_{3}\Gamma_{1}$. With the help of $\mathcal{R}$, we obtain the two-body states in the form
@@
where $\mathcal{N}_{J,M_{1};K,M_{2}}^{(4I-2n,0)}$ is the normalization coefficient that is given by $SO(5)$ isoscalar (see S.M.Sec.VI), and $f_{J_{1},M_{1};J_{2},M_{2}}^{4I-2n}$ is a symmetric homogeneous polynomial with the degree of 4I2n4I-2n defined as
@@
where $C_{j_{1},m_{1};j_{1}^{\prime},m_{1}^{\prime}}^{J_{1},M_{1}}$ is the SU(2) group Celebsh-Gordon coefficient . The statistics of the two-body wavefunciton is captured by the power of $S^{n}$: Odd and even values of $n$ correspond to the fermionic and bosonic statistics, respectively.
We move forward to the two-body states in more complex channels with $n_{1}>n_{2}$. Let us take the case of $n_{1}-n_{2}=1$ as an example which involves the regular $SO(5)$ harmonics in the absence of Yang monopole. In analogy to the SO(5) invariant $S$, an $SO(5)$ vector, defined as $X_{a}(x_{a},n_{i};x_{a}^{\prime},n_{i}^{\prime})=\psi_{\alpha}(x_{a},n_{i})(\mathcal{R}\Gamma_{a})_{\alpha\beta}\psi_{\beta}(x_{a}^{\prime},n_{i}^{\prime})$, transforms under the IRREP $(1,1)_{so5}$ . $X_{a}$ can be decomposed into an SO(4) scalar $X_{5}$ and a 4-vector $X_{1\sim 4}$ transforming as $(\frac{1}{2};\frac{1}{2})_{so4}$(see S.M.sup (2020).Table(S3)). $S$ and $X_{5}$ are orthogonal to each other, since the former is SO(5) invariant and the latter is a component of five-vector. According to group theoretical analysis, the SO(5) of IRREP $(4I-2n-1,1)_{so5}$ with $0\leq n\leq 2I-1$ can be decomposed into two branches of SO(4) IRREPs as follows,
@@
Hence, we have two sets of the two-body wavefunctions as
@@
in which $\Psi_{\frac{1}{2},\pm\frac{1}{2};\frac{1}{2},\pm\frac{1}{2}}^{(1,1)}=\frac{1}{\sqrt{2}}(X_{1}\pm iX_{2})$, and $\Psi_{\frac{1}{2},\pm\frac{1}{2};\frac{1}{2},\mp\frac{1}{2}}^{(1,1)}=\frac{1}{\sqrt{2}}(X_{3}\pm iX_{4}).$ More generally, for $(n_{1}+n_{2},n_{1}-n_{2})_{so(5)}$, the two-body wavefunctions in these irreps are obtained using higher-rank $SO(5)$ harmonics, as detailed in S. M. Sec(IV).
To implement numerical calculations, we generalize the pseudo-potential formalism Haldane (1983); Jain (2007) from $S^{2}$ to $S^{4}$. The pseudo-potential, based on the two-body states constructed above, is represented by a projection operator as follows
@@
The determinant-type Laughlin wavefunction takes the form of the Slater determinant raised to the $m$-th power.
@@
where $N$ is the particle number, $A_{h}$ is the abbreviation of $SO(4)$ IRREP $(j_{h},m_{h,1};k_{h},m_{h,2})$ for the $h$-th particle,and $P$ is permutation from 1 to $N$. The filling of Eq.(12) is $\nu=\frac{d(2I,0)}{d(2mI,0)}$ , which approaches $1/m^{3}$ in the large-$m$ limit.
@@
$\Psi^{(m=3)}_{\mbox{det},N=4}=\left(\varepsilon_{P_{1}P_{2}P_{3}P_{4}}\prod\limits_{j=1}^{4}\psi_{j}(x_{P_{j},a},n_{P_{j},i})\right)^{3}$ (13)
By counting the power of SO(5) spinors, the single-particle LLL orbitals involved in Eq.(13) correspond to the case of $mI=\frac{3}{2}$ with the degeneracy $d(3,0)=20$. The full-shell determinant state is $SO(5)$ invariant; its power, the Laughlin determinant state, is likewise an $SO(5)$ singlet. Upon exchanging two particles, the wavefunction acquires a phase factor $(-1)^{3}=-1$, which ensures Fermi statistics. Hence, when expanding Eq.(13), the relative wavefunction between a pair of particles $k$ and $l$ only contains
@@
Now we construct the pseudo-potential Hamiltonian for the many-body wavefunctions Eq.(12). The pseudo-potential Hamiltonian, constructed from two-body projection operators, is positive-definite and annihilates the wavefunction (12). Consequently, the annihilated wavefunction is a zero-energy ground state. For concreteness, the special case of Eq.(13) with $m=3,N=4$ is employed as an example, i.e.. Consider two fermions in the LLL orbits with $I^{\prime}=mI=\frac{3}{2}$. The two-body states satisfying the fermionic statistics are decomposed into SO(5) IRREPs of $(4I^{\prime}-2r-1-s,2r+1-s)_{so5}$ with $0\leq r<I^{\prime}$ and $0\leq s\leq 2r+1$. According to the two-body wavefunctions given in Eq.(VI.8) in S. M. Sec. (VI), the contraction number equals $2r+1$. Hence, the $r=0$ irreps, (4,0)SO(5)(4,0){SO(5)} and (5,1)SO(5)(5,1){SO(5)}, are orthogonal to the two-body sectors of the wavefunction in Eq. (13), as revealed by the contraction numbers. The above analysis can be easily generalize to an arbitary value of $N$ at $m=3$. Then the pseudo-potential Hamiltonian is constructed as
@@
where $P_{m,n}(k,l)$ is the projection operator of two-particle states of the $k$-th and $l$-th ones.
@@
Next we present the exact digonalization results for the case of $m=3$ and $N=4$. The $SO(5)$ is a rank-2 Lie group with two commutable generators, which can be taken as the sum of the diagonal operators of two inter-commutable $SU(2)$ subgroups defined in Eq.(4) of every particle: $J_{3}=\sum_{l=1}^{N}J_{3}(l)$ and $K_{3}=\sum_{l=1}^{N}K_{3}(l)$. The pseudo-potential Hamiltonian is diagonalized in blocks labeled by different values of $(M_{J},M_{K})$, which are $J_{3}$ and $K_{3}$ eigenvalues, respectively. Previously analysis shows that the wavefunction of Eq.(13) is an SO(5) singlet and a ground state of the Hamiltonian Eq. (15). Diagonalizations are performed and the ground state is found non-degenerate, hence, the wavefunction of Eq.(13) is the unique ground state. Energy spectra in the blocks of $(0,0)$ and $(\frac{1}{2},\frac{1}{2})$ with spectra shown in Fig.(1). The energies of the lowest and next lowest excited states are very close at $4.80$ and $4.93$ under the unit of $\frac{\hbar^{2}}{2MR^{2}}$. The lowest excited state is also an $SO(5)$ scalar as shown in the sector of $(0,0)$, and the next lowest excited states are 35-fold degenerate forming the SO(5) IRREP of $(4,0)_{{so5}}$. Such an $SO(5)$ IRREP is decomposed into different $SO(4)$ IRREPs carrying integer quantum numbers of $(2,0)_{so4}\oplus(1,1)_{so4}\oplus(0,2)$ with dimensions of $5$, 99, and $5$, respectively, and those carrying half-integer quantum numbers of $(\frac{3}{2},\frac{1}{2})_{so4}\oplus(\frac{1}{2},\frac{3}{2})_{so4}$ with dimensions 8 and 8, respectively, as shown in Fig.(1) . More spectra details are shown on S.M.sup (2020).Sec(VIII).
@@
We now present the spectra of excitations of the quasi-hole ($N=3$) and quasi-particle ($N=5$) sectors as shown in Fig.(2) ($a$) and ($b$), respectively. Since each SO(5) IRREP contains SO(4) IRREPS $(j_{1},j_{2})_{so4}\oplus(j_{2},j_{1})_{so4}$ symmetrically, we only present the state in the sector $(\frac{1}{2},0)$ in terms of $(M_{J},M_{K})$ as a representative. As for the quasi-hole states of $N=3$, its ground states remains at zero energy with 20-fold degeneracy, lying in the SO(5) IRREP of $(3,0)_{so5}$, which is decomposed to SO(4) IRREPs of $(\frac{3}{2},0)_{so4}\oplus(0,\frac{3}{2})_{so4}\oplus(1,\frac{1}{2})_{so4}\oplus(\frac{1}{2},1)_{so4}$. In contrast, as for the quasi-particle states with $N=5$, there exist a finite excitation gap. The first excited states form the SO(5) IRREP of $(3,0)_{so5}$, which is the same as the quasi-hole case.
@@
Early studies of FQHE debated whether the ground state should be an incompressible quantum liquid or a Wigner crystal Fukuyama et al. (1979); Yoshioka et al. (1983); Yoshioka and Lee (1983); Levesque et al. (1984); Yoshioka (1984). The pair distribution function is frequently employed to reveal the incompressible nature of quantum liquids Jain (2007). We calculate the probability $h(\theta)$ of two particles separated at a distance, which is a function of the angle $\theta$ between two unit vectors from the center to particles due to the rotation symmetry. It is defined as